inyoskin:

M is the midpoint ofJK. The coordinates of J are (6, 3) and the coordinates of M are (-3, 4), find the coordinates of K.

Shadow:

\[M = ( \frac{ x _{1} + x _{2}}{ 2 } , \frac{ y _{1} + y _{2} }{ 2 })\]

Shadow:

This is the formula for the midpoint, simple separate it into: \[x = \frac{ x _{1} + x _{2}}{ 2 } , y = \frac{ y _{1} + y _{2} }{ 2 }\\]\] Where x is the x coordinate of the midpoint, and x1 is the x coordinate of J and x2 is the x coordinate of K. The same goes for y. Input your variables then solve.

Shadow:

@inyoskin Does that make sense?

inyoskin:

I'm trying to understand it, thank you for the formula. i am gong to use it now.

Shadow:

No problem

inyoskin:

is the answer (-6,2.5) or (-12,5) ?

Shadow:

The second one

inyoskin:

Thanks again i'm understanding it a lot better now.